OpenStudy (anonymous):

Solve x2 − 3x = 28. HELP ME (:

OpenStudy (anonymous):

x^2 -3x=28 x^2-3x-28=0 factor (x )(x )=0 think of two number that add up to -3 and multiply to -28

OpenStudy (anonymous):

what two number multiples to be 28?

OpenStudy (anonymous):

ok, one second. @modphysnoob

OpenStudy (anonymous):

-7, and 4

OpenStudy (anonymous):

@modphysnoob

OpenStudy (anonymous):

(x-7)(x+4)=0

OpenStudy (anonymous):

good job

OpenStudy (anonymous):

yep, -7+4=-3 -7*4=-28

OpenStudy (anonymous):

haha thanks @nader1

OpenStudy (anonymous):

so you'd get here (x-7)(x+4)=0 from here , we can deduce that x=7,-4

OpenStudy (anonymous):

thanks doe